Grade 12th passAlgebra

pls answer this question
a:b=100:75
b:c=10050
a:c+?

Rathod Hetal Bhagat
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Dear Student,
a/b = 100/75 ------(1)
       b/c = 100/50 ------(2)
       multiply (1) and (2) to get
       a/c = 8/3
       therefore a:c = 8:3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)